OpenStudy (anonymous):

I am confused the question was given to me as so.. 1/10 divided by 6/19 the choices are A. 19/60 B. 79/190 C. 3/95 D. (3) 3/19 I don't know if they flipped it (reciprocal) so i got two answers A and C. Help?

OpenStudy (godlovesme):

\[1/10 \div 6/19 = 1/10 \times 19/6\]

OpenStudy (anonymous):

yes. you flip the 2nd one and multipy instead of divide

OpenStudy (godlovesme):

yep :)

OpenStudy (anonymous):

thank you!

OpenStudy (godlovesme):

anytime :)

OpenStudy (godlovesme):

thanks for the medal guys :)

OpenStudy (anonymous):

Wait i gott (1) 4/15

OpenStudy (godlovesme):

did u change it to mixed #?

OpenStudy (anonymous):

OHHH

OpenStudy (anonymous):

ITS A

OpenStudy (anonymous):

i simplified wrong, my teacher tried to teach us a tip but i should not use it

OpenStudy (anonymous):

yeah its 19/60

OpenStudy (godlovesme):

yes

OpenStudy (anonymous):

i knew it was either a or c

OpenStudy (anonymous):

thank you!

OpenStudy (godlovesme):

good & my pleasure